OpenStudy (anonymous):

Write an expression for the general or nth term, an, for the geometric sequence. 1, -4, 16, -64, . . . A. an = 1 • (-4) n-1 B. an = 1 • (-4)n C. an = 1 • 4n-1 D. an = -1 • 4n

OpenStudy (anonymous):

the n , n-1 at the end is a power of 4,-4 right ?

OpenStudy (anonymous):

answer is B :) lol

OpenStudy (anonymous):

its not B ..

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

in order to find the solution you have to substitute n=1,2 in the equations and see what fit

OpenStudy (anonymous):

its b. plug 1 into A. you get -5. it isnt a power. that would be ^1

OpenStudy (anonymous):

michael0195 n-1 is a power of (-4)

OpenStudy (anonymous):

no its multiplying :) to make a power it would be (-4) ^n-1 there is no "^" its not a power :)

OpenStudy (anonymous):

if its a multiplication then non of the answers is correct.. plug n =2 into your favorite b and get -8

OpenStudy (anonymous):

there is no 2 in the sequesnt the sequence is 1, -4, 16, -64. plug those numbers into B and oyu get the next number in the sequence :)

OpenStudy (anonymous):

sequence*

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

but the numbers you plug for n are 1,2,3,4,.....

OpenStudy (anonymous):

NO! its A

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

why would they be 1,2,3,4 when it gives you the sequence in the question?

OpenStudy (anonymous):

mollyann please tell us if for exmaple in A the n-1 is a power of (-4) ?

OpenStudy (anonymous):

\[A. an = 1 * (-4)^{n-1}\]

OpenStudy (anonymous):

does A looks like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A is the correct answer

OpenStudy (anonymous):

and michael .. n represents natural number .. any natural number so n =1,2,3,4....

OpenStudy (anonymous):

for n =1 you should get 1 n = 2 you should get -4 and so ..

OpenStudy (anonymous):

in the future if there is superscript thats like raised as an exponent, you gotta type the carrot top "^" to tell us it is! lol

OpenStudy (anonymous):

whatever .-.

OpenStudy (anonymous):

or answer these question i cant figure them out :/ and i need A on this test or i fail ):

OpenStudy (anonymous):

Express the repeating decimal number as a quotient of two integers. 28.81 A . 3125 over 11 B. 83 over 11 C. 317 over 11 D. 5 over 52

OpenStudy (anonymous):

x = 0.81 (repeating) so 100x = 81.81 (repeating) and so 100x -x = 81.81 - 0.81 (repeating) 100x -x = 81 so: 99x = 81 x = 81/99 x= 9/11 28 + (9/11) = (28 * 11 + 9) / 11 = (308 +9) / 11 = 317 / 11

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw